16 yers later,Usha will be two times as old as she was 16 years ago.How old is she now.
Answer by oscargut(2103)

Hi,
Solution:
Let x = age of usha now
x+16 = 2(x-16)
x+16 = 2x-32
-x = -48
x = 48
Answer: 48 years

If you have any doubt or if you have more problems my e-mail is:
mthman@gmail.com